# D4C glove-train: solving the RPM and Bongard-logo problem by distributing and Circumscribing concepts

### 17 Mar 2024 v3

@@ VIII-B Experiment on Bongard-Logo
Consequently, we conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN as backbones for D2C, D3C, and D4C. The experimental results are presented in the table III. In the experiment with D3C, we conducted an additional experiment on its lightweight version, D3C-cos, to verify that replacing D3C with D3C-cos would not significantly affect performance. As can be seen in the table III, we did not conduct an experiment with Triple-CFN+D3C due to the large model size of Triple-CFN. This made it impractical to deploy the model using conceptual distribution and logical optimization with the Sinkhorn distance on our existing experimental equipment. This was the driving force behind proposing D3C-cos as an alternative to D3C in this paper. In addition, the combination of D4C and PMoC+Straw Pose-Transformer has not been implemented due to limitations in computing power.
@@ IX Conclusion
In summary, D2C presents a novel approach to tackling Bongard-Logo and RPM problems by resampling training data and introducing new loss function terms, effectively outlining new boundaries for solutions to these challenges. However, the extent to which these boundaries adapt to the specific problems remains uncertain. Conversely, D3C advocates for representing concepts in abstract reasoning problems using distributions. It posits that, compared to vectors, distributions are better suited to capture the high-dimensionality, abstractness, and complexity inherent in human concepts, and experimental evidence underscores the superiority of this approach.

### 25 Mar 2024 v4

@@ VIII-B Experiment on Bongard-Logo
Consequently, we conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN as backbones for D2C, D3C, and D4C and using PMoC as backbone for D2C and D4C. The experimental results are presented in the table III. In the experiment with D3C, we conducted an additional experiment on its lightweight version, D3C-cos, to verify that replacing D3C with D3C-cos would not significantly affect performance. As can be seen in the table III, we did not conduct an experiment with Triple-CFN+D3C due to the large model size of Triple-CFN. This made it impractical to deploy the model using conceptual distribution and logical optimization with the Sinkhorn distance on our existing experimental equipment. This was the driving force behind proposing D3C-cos as an alternative to D3C in this paper. In addition, the combination of D4C and PMoC+Straw Pose-Transformer has not been implemented due to limitations in computing power.
@@ IX Conclusion
In summary, D3C advocates for representing concepts in abstract reasoning problems using distributions. It posits that, compared to vectors, distributions are better suited to capture the high-dimensionality, abstractness, and complexity inherent in human concepts, and experimental evidence underscores the superiority of this approach. Additionally, D2C presents a novel approach to tackling Bongard-Logo and RPM problems by resampling training data and introducing new loss function terms, effectively outlining new boundaries for solutions to these challenges. However, the extent to which these boundaries adapt to the specific problems remains uncertain.